Calculating and Reporting Deferred Income Taxes

Bens' Corporation paid $15,000 on December 31, 2016, for equipment with a three-year useful life. The equipment will be depreciated in the amount of $5,000 each year. Bens' took the entire $15,000 as an expense in its tax return in 2016. Assume this is the only timing difference between the firm's books and its tax return. Bens' tax rate is 40%.

Required
a. What amount of deferred tax liability should appear in Bens' 12/31/2016 balance sheet?
$Answer

b. Where in the balance sheet should the deferred tax liability appear?

Deferred tax liabilities are recorded as noncurrent liabilities on the balance sheet.

Deferred tax liabilities are recorded as current liabilities on the balance sheet.

Deferred tax liabilities are recorded as current and/or noncurrent liabilities on the balance sheet, depending on when they are due.

Deferred tax liabilities are not reported on the balance sheet, rather they are disclosed in the Notes to the Financial Statements.

c. What amount of deferred tax liability should appear in Bens' 12/31/2017 balance sheet?

Deferred tax liability Answer
Answer #1

Solution a:

Amount of deferred tax liability should appear in Bens' 12/31/2016 balance sheet = $15,000*40% = $6,000

Solution b:

Deferred tax liabilities are recorded as current and/or noncurrent liabilities on the balance sheet, depending on when they are due.

Hence 3rd option is correct.

Solution c:

mount of deferred tax liability should appear in Bens' 12/31/2017 balance sheet = ($15,000 - $5,000)*40% = $4,000